OpenStudy (anonymous):

Find the area of the plane region that is enclosed by the curves: \[y=\left| x^2-4 \right|\] and\[y=(\frac{ x^2 }{ 2 }) +4\]

OpenStudy (anonymous):

I just need to know what to do with the absolute value sign.

OpenStudy (anonymous):

@SmoothMath @Zarkon @UnkleRhaukus

OpenStudy (anonymous):

@Hero

hero (hero):

You graph it. Remember y = |x^2 -4| = x^2 - 4 or 4 - x^2

OpenStudy (anonymous):

So it's really 3 curves I am dealing with right?

hero (hero):

You have to pay attention to the OR part. You will have two separate graphs. However, my theory is that the areas of both will be the same.

OpenStudy (anonymous):

Okay now I am confused.

OpenStudy (anonymous):

I get that i consider the or part but then how do I find the area?

hero (hero):

On one graph, you will find the area enclosing y = x^2 - 4 and y = x^2/2 + 4 On the other graph you will find the are enclosing y = 4 - x^2 and y = x^2/2 + 4

OpenStudy (anonymous):

Ohh okay! So I find the area between the curves for both cases and add them up right?

hero (hero):

In theory

OpenStudy (anonymous):

Hmm let me do it...

hero (hero):

Don't forget to find the enclosing points first.

OpenStudy (anonymous):

Yeah that's easy.

OpenStudy (anonymous):

yeah I know that @UnkleRhaukus but I jest need to know how to deal with the absolute value sign.

OpenStudy (unklerhaukus):